# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v5

@@ 6 Experimental Methodology
We extend Ramulator [72, 128] to implement and evaluate the five SMD maintenance mechanisms (SMD-FR, SMD-VR, SMD-PRP, SMD-PRP+, and SMD-MS) that we describe in section 5. We use DRAMPower [2, 14] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [88]. We warm-up the caches by fast-forwarding 100 million instructions. We simulate each representative trace for 500 million instructions (for multi-core simulations, we simulate until each core executes at least 500 million instructions). We release the source code of modified version of Ramulator openly and freely available at [133].
@@
| Processor |
& 4-wide issue CPU core, 1-4 cores, 8 MSHRs/core, 128-entry instruction window |
| Last-Level Cache |
cache-line, 8-way associative, /core |
| Memory Controller |
64-entry read/write request queue, FR-FCFS-Cap [104] |
| DRAM |
DDR4-3200 [50], refresh period, 4 channels, 2 ranks, 4/4 bank groups/banks, 128K-row bank, 512-row subarray, row size |
Workloads. We evaluate 44 single-core applications from four benchmark suites: SPEC CPU2006 [144], TPC [147], STREAM [93], and MediaBench [32]. We classify the workloads in three groups based on their memory intensity, which we measure using misses-per-kilo-instructions (MPKI) in the last-level cache (LLC). The low, medium, and high memory intensity groups consist of workloads with , , and . We randomly combine single-core workloads to create multi-programmed workloads. Each multi-programmed workload group, 4c-low, 4c-medium, and 4c-high, contains four-core workloads with low, medium, and high memory intensity.
@@ 7.1 Single-core Performance
Fig. 10 shows the speedup of single-core workloads when using the following maintenance mechanisms: 1) SMD-FR, 2) SMD-VR, 3) SMD-PRP, 4) SMD-PRP+, 5) SMD-MS, and 6) the combination of SMD-VR, SMD-PRP, and SMD-MS (SMD-Combined).101010We individually evaluate SMD-PRP, SMD-PRP+, and SMD-MS using SMD-FR as a refresh mechanism since SMD-FR is very simple and easy to implement. In SMD-Combined, we prefer SMD-PRP over SMD-PRP+ to keep the DRAM chip area overhead minimal. Based on [86], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a -bit Bloom Filter with hash functions. SMD-PRP and SMD-PRP+ refresh the victims of an activated row with a high probability, i.e., . SMD-MS operates with an aggressive 5-minute scrubbing period. 111111We analyze SMD-PRP, SMD-PRP+, and SMD-MS at various configurations and find that their overheads are relatively low even at more aggressive settings. We omit the details due to space limitations. We calculate the speedup compared to DDR4 DRAM with a nominal refresh period. To demonstrate the maximum speedup achievable by eliminating the entire DRAM refresh overhead, Fig. 10 also includes the performance of hypothetical DRAM that does not require refresh (NoRefresh). The figure shows workloads that have at least 10 Last-Level Cache Misses per Kilo Instruction (MPKI). GMEAN-22 provides the geometric mean of the workloads and GMEAN-ALL of all single-core workloads that we evaluate.
We make five key observations. First, SMD-FR provides 5.7% average speedup (up to 23.2%) over conventional DDR4 refresh (84.1% of the speedup that NoRefresh provides). Although SMD-FR and DDR4 have similar average latency to refresh a single DRAM row, SMD-FR outperforms DDR4 refresh due to two main reasons: 1) enabling concurrent access and refresh of two different lock regions within a DRAM bank and 2) completely eliminating the REF commands on the DRAM commands bus to reduce the command bus utilization. SMD-FR requires very low area overhead and no information with respect to DRAM cell characteristics, yet provides significant speedup.
Second, SMD-VR provides 6.1% average speedup (up to 24.5%), achieving 89.9% of the speedup that NoRefresh provides. SMD-VR provides higher speedup due to eliminating unnecessary refresh operations to retention-strong rows. Compared to SMD-FR, SMD-VR’s performance benefits are limited because SMD-FR significantly mitigates the refresh overhead by overlapping refresh operations with accesses to other lock regions, leaving small opportunity for further improvement by skipping refreshes to retention-strong rows.
Third, both SMD-FR and SMD-VR perform close to the hypothetical “NoRefresh” DRAM. On average, SMD-FR/SMD-VR provide 83.8%/86.8% of the speedup that “NoRefresh” provides compared to the DDR4 baseline.
Fourth, although the overheads of the SMD-PRP, SMD-PRP+, and SMD-MS mechanisms partially negate the performance benefits of SMD-FR, they still achieve average (maximum) speedup of 4.7%/5.7%/5.4% (9.4%/23.2%/12.9) over the DDR4 baseline, when integrated on top of SMD-FR.
Fifth, SMD-Combined improves performance by 4.8% (up to 10.6%), showing that SMD enables robust RowHammer protection and frequent (5-minute) memory scrubbing while still improving system performance.
@@ 7.2 Multi-core Performance
Fig. 11 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) for 60 four-core workloads (20 per memory intensity level). The bars (error lines) represent the average (minimum and maximum) weighted speedup across the 20 workloads in the corresponding group. We make three major observations.
First, we observe the largest average speedup in memory-intensive four-core workloads, i.e., 4c-high. SMD-FR/SMD-VR improve average speedup by 8.7%/9.2% (9.9%/10.2 maximum), which corresponds to 86.8%/91.2% of the speedup that hypothetical “NoRefresh” achieves over DDR4.
Second, combining SMD-VR+SMD-PRP+SMD-MS provides 7.6% average (9.1% maximum) improvement in weighted speedup, which indicates that SMD-PRP and SMD-MS incur only small performance overhead on top of SMD-VR.
Third, the performance benefits become smaller as the memory intensity of the workloads decrease. Specifically, 4c-medium achieves lower speedup than 4c-high, and 4c-low achieves lower speedup than 4c-medium.
Performance Summary. The new DRAM refresh mechanisms SMD-FR and SMD-VR significantly improve system performance, achieving speedup close to that provided by the hypothetical “NoRefresh” DRAM. SMD enables efficient and robust RowHammer protection (SMD-PRP and SMD-PRP+) and memory scrubbing (SMD-MS) that still provide significant speedup when integrated with SMD-FR. Overall, SMD enables DRAM that is both faster and more robust than DDR4.
7.3 Energy Consumption
@@
Fig. 12 shows the average DRAM energy consumption normalized to the DDR4 baseline for single- and four-core workloads. We make three major observations.
First, SMD-FR reduces DRAM energy for both single- and four-core workloads (by 5.9% and 5.5% on average) over DDR4 mainly due to reduction in execution time.
Second, SMD-VR reduces DRAM energy by 6.8%/6.7% on average for single-core and 4c-high workloads by especially eliminating some unnecessary refresh operations. The maximum DRAM energy reduction that can be achieved by completely eliminating DRAM refresh operation is 7.8% as indicated by “NoRefresh” DRAM.
Third, SMD-Combined reduce DRAM energy by 5.2% for both single-core and 4c-high workloads. This shows that the energy overhead of SMD-PRP and SMD-MS are small.
The new SMD-based refresh mechanisms eliminate most of the energy overhead of DRAM refresh. SMD-based RowHammer protection and memory scrubbing improve DRAM reliability and security with low DRAM energy cost.
7.4 Isolating SMD’s Performance Benefits
We evaluate SMD’s performance without the added benefits of memory accesses and maintenance operation parallelization (i.e., SMD with a single lock region) using all four-core heavily memory-intensive (4c-high) workloads. This configuration of SMD, called SMD-Combined-1LR, implements the same maintenance mechanisms as SMD-Combined does. We compare SMD-Combined-LR to SMD-Combined. We also evaluate a DRAM chip that leverages refresh-access parallelization [16, 161], called DSARP. A system that implements DSARP can parallelize refresh and memory access operations to different subarrays. We compare DSARP to SMD-FR. Figure 13 shows the speedup of 4-core workloads using these three mechanisms (SMD-FR, DSARP, SMD-Combined-1LR) and NoRefresh.
We make two major observations from Figure 13. First, SMD-Combined-1LR improves performance by 0.9% (3.2%) on average (at maximum) across all workloads. SMD-Combined-1LR improves performance less than (and sometimes degrades performance) SMD-Combined because it cannot concurrently perform maintenance operations and memory accesses. Second, DSARP improves performance by 5.0% (6.0%) on average (at maximum) across all workloads. We conclude that being able to concurrently execute maintenance and memory access operations at the same time increases SMD’s performance benefits, and SMD-FR provides comparable performance benefits to a DRAM chip that can leverage refresh-access parallelization (DSARP).
@@ 7.5.1 DRAM Refresh Period
7.5.2 Lock Region Size
The size of a SMD lock region affects the performance and energy of SMD-based maintenance mechanisms. Generally, it is desirable to have many small-sized lock regions to enable locking smaller portions of the DRAM chip for maintenance and allow accesses to non-locked regions. However, the DRAM chip area overhead of SMD increases as the number of lock regions in a bank increase. We analyze the performance and energy impact of different lock region size configurations.
We use the following equation to calculate the number of lock regions per bank () by dividing the number of rows in a bank () to the number of rows in a lock region ().
Figure 15 shows the speedup and DRAM energy savings that SMD-FR, SMD-VR, and SMD-Combined achieve for different number of lock regions per bank across 4c-high workloads. We make two key observations. First, with a single lock region per bank, SMD-FR/SMD-Combined on average perform 3.7%/8.9% worse than the DDR4 baseline, which leads to 3.2%/10.7% higher average DRAM energy consumption. This is because a maintenance operation locks the entire bank, and thus the MC waits for ARI upon receiving an ACT_NACK to issue another ACT to the same bank. Although SMD-VR suffers from the same issue, it still outperforms the baseline by eliminating unnecessary refreshes to rows with high retention times.
Second, when the lock regions are smaller, and thus a bank has four or more lock regions, all mechanisms outperform the baseline for all 4c-high workloads and save DRAM energy. This shows that even a small number of lock regions are sufficient to overlap the maintenance operation latency with accesses to non-locked regions.
Third, the performance and energy benefits of increasing the number of lock regions gradually diminish. We find that 8192-row lock regions incur low area overhead and achieve a significant portion of the speedup that smaller lock regions provide (e.g., SMD-Combined achieves 7.7%/9.3% speedup with 16/256 lock regions). Therefore, by default, we use lock regions per bank in our evaluations.
We conclude that SMD significantly reduces the overhead of DRAM maintenance operations even with large lock regions.
@@ 7.5.8 SMD’s Performance Sensitivity to Number of Cores
7.6 Hardware Overhead
1) alert_n. alert_n is currently used to inform the MC that the DRAM chip detected a CRC or parity check failure on the issued command, and thus the MC must issue the command again. The alert_n signal can simply be asserted not only on CRC or parity check failure, but also when the MC attempts accessing a row in a locked region. This approach does not require an additional physical pin. However, assigning multiple meanings to an already ambiguous alert_n signal [50] could complicate MC design.
2) Introducing a new pin. To potentially simplify MC design, SMD can introduce a new pin to transmit ACT_NACK signals. Only one new pin between the MC and all memory devices (e.g., all DRAM chips in a memory channel or a rank) is enough for systems that use rank-based DRAM chip organizations (e.g., DDR4 DIMMs). For example, 128 new pins are needed for a high-end system with 32 memory channels equipped with 4-rank memory chips in each channel. However, a system at this scale already has 4’094 processor pins [153], and 128 new pins would amount to a relatively small 3.13% increase in pins. Similar to how alert_n signal works, the module can send the MC a single ACT_NACK when any of the per-chip ACT_NACK signals are asserted.
7.6.1 DRAM Chip Modifications
We use CACTI [100] to evaluate the hardware overhead of the changes that SMD introduces over a conventional DRAM bank (highlighted in Fig. 3) assuming technology. Lock Region Table (LRT) is a small table that stores a single bit for each lock region to indicate whether or not the lock region is under maintenance. In our evaluation, we assume that a bank is divided into lock regions. Therefore, LRT consists of only bits, which are indexed using a 4-bit lock region address. According to our evaluation, an LRT incurs only area overhead per bank. The area overhead of all LRTs in a DRAM chip is only 0.001% of a DRAM chip. The access time of an LRT is , which is only 0.4% of typical row activation latency (tRCD) of .
SMD adds a per-region RA-latch to enable accessing one lock region while another is under maintenance. An RA-latch stores a pre-decoded row address, which is provided by the global row address decoder, and drives the local row decoders where the row address is fully decoded. According to our evaluation, all RA-latches incur a total area overhead of 1.63% of a DRAM chip. An RA-latch has only latency, which is negligible compared to tRCD.
Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the maintenance mechanisms presented in section 5. The simple refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. SMD-VR maintains bloom filters that constitute a large portion of the required . The DRAM chip area overhead of each refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-PRP requires area, which is a similarly small portion of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. Across a DRAM chip, SMD-DRP requires area, which is 7.0% of a typical DRAM chip size. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
7.6.2 Memory Controller Modifications
We slightly modify the MC’s request scheduling mechanism to retry a rejected ACT command as we explain in section 4.2. Upon receiving an ACT_NACK, the MC marks the corresponding bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FRFCFS [104]).
No further changes are required in the MC to support different maintenance mechanisms enabled by SMD. Thus, SMD enables the DRAM vendors to update existing maintenance mechanisms and implement new ones without any further changes to MCs that support SMD.

### v6

@@ 6 Experimental Methodology
We extend Ramulator [82, 81] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section 5. We use DRAMPower [112, 113] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [114]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
@@
Processor | & 4-wide issue CPU core, 1-4 cores, 8 MSHRs/core, 128-entry instruction window |
Last-Level Cache | cache-line, 8-way associative, /core |
Memory Controller | 64-entry read/write request queue, FR-FCFS-Cap [115] |
DRAM | DDR4-3200 [78], refresh period, 4 channels, 2 ranks, 4/4 bank groups/banks, 128K-row bank, 512-row subarray, row size |
DDR4 Baseline. The baseline system 1) uses per-rank refresh111111The DDR4 standard does not support per-bank refresh. and 2) does not perform ECC-scrubbing.
Workloads. We evaluate 62 single-core applications from four benchmark suites: SPEC CPU2006 [116], SPEC CPU2017 [117], TPC [118], STREAM [119], and MediaBench [120]. We classify the workloads in three memory intensity groups measured using misses-per-kilo-instructions (MPKI) in the last-level cache (LLC): low (), medium (), and high (). We randomly combine single-core workloads to create multi-programmed workloads. Each multi-programmed workload group, 4c-low, 4c-medium, and 4c-high, contains four-core workloads.
Metrics. We use Instructions Per Cycle (IPC) to evaluate the performance of single-core workloads. For multi-core workloads, we evaluate the system throughput using the weighted speedup metric [121, 122, 123].
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [16] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode to reduce delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [16] implements DARP and modifies DRAM chips and the interface to perform a refresh operation in one subarray and a memory accesses in another subarray and thereby hide the latency of refresh operations and reduce delays they impose on memory requests. We modify DARP and DSARP such that they consider victim row refresh operations issued by Deterministic RowHammer Protection (section 5.2) and attempt to reduce delays these operations impose on memory requests.
Configuring Maintenance Mechanisms. SMD-FR refreshes a DRAM row every . Based on [17], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a -bit Bloom Filter with hash functions. SMD-PRP and SMD-PRP+ refresh the victims of an activated row with a high probability, i.e., . SMD-DRP refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS operates with an aggressive 5-minute scrubbing period.
Evaluated System Configurations. SMD-FR-1LR implements a single lock region in a bank, i.e., it cannot concurrently perform a refresh operation and a memory access. DSARP/SMD-FR can concurrently perform a refresh in a subarray/lock region and a memory access in another. MC-Combined-DARP, MC-Combined-DSARP, and SMD-Combined implement the memory-controller-(MC) and SMD-based versions of Fixed Rate Refresh (section 5.1), Deterministic RowHammer Protection (section 5.2), and Memory Scrubbing (section 5.3). MC-based implementations of SMD maintenance mechanisms are functionally equivalent to SMD-based ones (e.g., SMD-DRP and MC-DRP are configured to provide the same RowHammer prevention guarantees).
@@ 7.1 Single-core Performance
Fig. 9 shows the speedup of the 22 highest MPKI single-core workloads for DARP, DSARP, SMD, and a hypothetical No-Refresh configuration that does not perform any maintenance operations, over the Baseline system.
We make three key observations. First, SMD-Combined provides a comparable average speedup of 5.9% to No-Refresh’s 5.0% as SMD-Combined eliminates a large fraction of the maintenance overhead by allowing the memory controller to access lock regions that are not under maintenance. Second, SMD-Combined outperforms MC-Combined-DSARP and MC-Combined-DARP by 0.6% and 3.9%, respectively. While MC-Combined-DSARP can concurrently access main memory and perform a maintenance operation, it needs to issue a DRAM command (e.g., a per bank refresh command for periodic refresh) for each maintenance operation and incur delays for other DRAM commands that the memory controller issues to serve main memory requests. In contrast, SMD-Combined autonomously performs each maintenance operation inside the DRAM chip and does not incur delays for DRAM commands (except those incurred by ACT_NACKs). MC-Combined-DARP cannot concurrently perform a maintenance operation and access main memory, thereby performs worse than both SMD-Combined and MC-Combined-DSARP. We make similar observations on SMD-FR’s, DSARP’s, and DARP’s performance. These observations are in line with prior work [16]. Third, intelligent refresh scheduling (DARP) provides 1.1% average speedup across all workloads by reducing the impact of refresh operations on the latency of memory requests. SMD-FR-1LR incurs 2.9% slowdown over DARP. Although SMD-FR-1LR locks a bank only when there is no active row in a bank, it does not have a global view over memory requests in memory controller read/write queues, unlike DARP, and cannot opportunistically perform refresh operations while the memory controller is in write mode.
@@ 7.2 Multi-core Performance
Fig. 10 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) for 60 four-core workloads (20 per memory intensity level). The white circles inside each box (error lines) represent the average (minimum and maximum) weighted speedup across the 20 workloads in the corresponding group.
We make three major observations from Fig. 10. First, SMD-Combined provides 8.6% and 4.1% speedups on average across 4c-high workloads over MC-Combined-DARP and MC-Combined-DSARP, respectively. We attribute these speedups to i) SMD’s maintenance-access parallelization and ii) SMD’s ability to perform maintenance operations autonomously inside the DRAM chip without the memory controller having to issue DRAM commands. Second, SMD-FR-1LR incurs an average 4.5% slowdown on average across 4c-high workloads compared to DARP. While DARP can reduce the overheads of periodic refresh by intelligently scheduling refresh operations, SMD-FR-1LR cannot. We attribute SMD-FR-1LR’s overheads to the relatively high rate of ACT_NACK commands (not shown in the figure) it issues: SMD-FR-1LR issues an ACT_NACK command for every 11.7 activate commands, on average across all 4c-high workloads. Third, we observe that i) the speedup and slowdown trends for the tested mechanisms are similar for the average 4c-medium workload, and ii) the tested mechanisms provide little performance improvement for the average 4c-low workload with very low memory intensity.
We conclude that SMD i) provides substantial system performance benefits by concurrently performing a maintenance operation and a memory access in different lock regions in a bank and ii) does not significantly hurt system performance when SMD performs maintenance at bank granularity (i.e., implements one lock region per DRAM bank).
7.3 SMD’s Energy Consumption
@@
Fig. 11 shows the average DRAM energy consumption normalized to the DDR4 baseline for single- and four-core workloads. We make two major observations.
First, SMD-Combined and SMD-FR reduce DRAM energy by 3.9% (4.3%) and 5.2% (5.2%) across all single-core (4c-high) workloads compared to the baseline, respectively. SMD-Combined performs close to the hypothetical No-Refresh configuration, providing 59.7% of the energy reduction benefits of No-Refresh on average across 4c-high workloads. We attribute the reduction in DRAM energy to i) the reduced DRAM background energy consumption because SMD shortens the execution time for these workloads and ii) eliminating DRAM commands for maintenance operations (e.g., for periodic refresh) on the power-hungry DDRx bus. Second, SMD-FR-1LR increases average single-core (4c-high) workload DRAM energy consumption by 0.1% (3.0%). While SMD-FR-1LR also eliminates DRAM commands for maintenance operations, it induces energy overheads mainly because of the increased execution time for the evaluated workloads.
7.4 Hardware Overhead
1) alert_n. alert_n is currently used to inform the MC that the DRAM chip detected a CRC or parity check failure on the issued command, and thus the MC must issue the command again. The alert_n signal can simply be asserted not only on CRC or parity check failure, but also when the MC attempts accessing a row in a locked region. This approach does not require an additional physical pin. However, 1) assigning multiple meanings to an already ambiguous alert_n signal [78] could complicate MC design and 2) alert_n, as currently defined in the standard [78], is an open drain signal that is slow and not precise enough.
2) Introducing a new pin. To potentially simplify MC design, SMD can introduce a new pin to transmit ACT_NACK signals. Only one new pin between the MC and all memory devices (e.g., all DRAM chips in a memory channel or a rank) is enough for systems that use rank-based DRAM chip organizations (e.g., DDR4 DIMMs). For example, 128 new pins are needed for a high-end system with 32 memory channels equipped with 4-rank memory chips in each channel. However, a system of this scale already has 4’094 processor pins [125], and 128 new pins would amount to a relatively small 3.13% increase in pins. Similar to how alert_n signal works, the module can send the MC a single ACT_NACK when any of the per-chip ACT_NACK signals are asserted.
DRAM Chip Modifications. We use CACTI [126] to evaluate the hardware overhead of the changes that SMD introduces over a conventional DRAM bank (highlighted in Fig. 2) assuming technology.121212Capitalizing on the latest DRAM technology libraries to implement the changes that SMD introduces would likely provide more accurate area overhead estimations. However, such libraries and tools are proprietary and we do not have access to them. Lock Region Table (LRT) is a small table that stores a single bit for each lock region in a DRAM bank to indicate whether or not the lock region is under maintenance. In our evaluation, we assume that a bank is divided into lock regions. Therefore, LRT consists of only bits, which are indexed using a 4-bit lock region address. According to our evaluation, an LRT incurs only area overhead per bank. The area overhead of all LRTs in a DRAM chip is only 0.001% of a DRAM chip. The access time of an LRT is , which is only 0.4% of typical row activation latency (tRCD) of .
SMD adds a per-region RA-latch to enable accessing one lock region while another is under maintenance. An RA-latch stores a pre-decoded row address, which is provided by the global row address decoder, and drives the local row decoders where the row address is fully decoded. This design builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. According to our evaluation, all RA-latches incur a total area overhead of 1.6% of a DRAM chip. An RA-latch has only latency, which is negligible compared to tRCD.
Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the maintenance mechanisms presented in section 5. The simple refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. Across a DRAM chip, SMD-DRP requires area, which is 7.0% of a typical DRAM chip size. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
Memory Controller Modifications. We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section 4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FRFCFS [115]).
Tracking locked regions. To apply the ARI timing parameter (e.g., to an ACT command targeting a locked region), the memory controller tracks which regions are locked. The memory controller could store the address of the locked region in every DRAM bank (since only one region can be under maintenance at a time in a bank). 1 rank address bit, 4 bank address bits, and 4 lock region address bits (9 bits) are sufficient to track every under-maintenance lock region in 2 ranks, 8 chips, and 16 banks (256 unique banks). Therefore, the storage cost for tracking locked regions is only 288 bytes for the evaluated memory channel with a dual-rank x8 memory module.
Address mapping schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section 7, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across DRAM channels (from the most significant to the least significant bit, the physical address is mapped to rows, banks, ranks, columns, and channels). We also evaluate SMD using the mapping scheme described in [127] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
No further changes are required in the MC to support different maintenance mechanisms enabled by SMD. Thus, SMD enables DRAM designers to update existing maintenance mechanisms and implement new ones without any further changes to MCs that support SMD.
@@ 7.5.7 SMD’s Performance Sensitivity to Number of Cores
8 Discussion
8.1 Prioritizing Memory Requests
To minimize the impact of rejected activate commands on system performance, an SMD chip could apply a different activate command rejection policy that prioritizes activate commands over maintenance operations. We describe and estimate the performance impacts of one such policy.
“Pause Maintenenace” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the lock region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section 5.1) during a periodic refresh maintenance operation. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Latency of Pausing Maintenance Operations. Even if the SMD chip decides to pause the ongoing maintenance operation, the activate command cannot immediately proceed because the lock region might not be in the precharged state (i.e., the DRAM bank might not be ready to activate a row). The DRAM chip requires up to (section 2.1) to bring the under-maintenance lock region’s state to the precharged state (in case an activate command was just issued by the maintenance mechanism), depending on the lock region’s state when the activate command is received.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR as SMD-FR that refreshes one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 18 shows weighted speedup for 60 four-core workloads (20 per memory intensity level) on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right.
We make two key observations from Fig. 18. First, SMD-PMP-FR provides a larger average speedup than SMD-FR for 4c-high workloads. SMD-PMP-FR’s ACT_NACK rate is 2% smaller than SMD-FR’s, which explains the 0.4% larger average speedup it provides over SMD-FR. Second, SMD-FR-1LR induces a smaller average system slowdown than SMD-PMP-FR-1LR. We conclude that an SMD-PMP design (whose performance could be accurately estimated using our methodology) that pauses maintenance operations to serve activate commands can improve performance.